OpenStudy (anonymous):

Oliver and Marie are graphing two equations on a coordinate grid. Oliver has graphed the equation y = 2x + 2. If Marie graphs y = 2x - 3, where will the y-intercept of her graph be in relation to Oliver's graph? A) 3 units lower B) 5 units lower C) 5 units higher D) the same place

OpenStudy (anonymous):

Oliver and Marie are graphing two equations on a coordinate grid. Oliver has graphed the equation y = 2^x + 2. If Marie graphs y = 2^x - 3, where will the y-intercept of her graph be in relation to Oliver's graph? A) 3 units lower B) 5 units lower C) 5 units higher D) the same place